OpenStudy (anonymous):

2(6(2/5)+7(1/3) ) i have the decimal answer but im supposed to present answer as a mixed number, integer, or proper fraction and my mind is shutting down lol

Directrix (directrix):

2 [ 6 (2/5) + 7 (1/3) ] --> Is this the problem? My eyes are shutting down.

OpenStudy (anonymous):

yes lol

OpenStudy (anonymous):

the numbers in parenthesis are mixed number fractions

Directrix (directrix):

What is your decimal answer? I want to see if my fractional answer is anywhere near it.

OpenStudy (anonymous):

rounded its 9.5 but i could be wrong

Directrix (directrix):

Are you saying that this --> 7 (1/3) is a mixed number?

OpenStudy (anonymous):

yes and the 6(2/5) is one also

Directrix (directrix):

Now, a mixed number is something like 2 and 3/4 which is written as 2 3/4 with no multiplication symbol or grouping symbol. 2 3/4 = 11/4 as an improper fraction. What is odd is that my fractional answer is close to your decimal answer but I did not see any mixed numbers.

OpenStudy (anonymous):

yea the equation here is written right but i judt didnt know how to type it right on here sorry its 2(6 2/5+7 1/3)

Directrix (directrix):

The problem is then: 2 [ 6 2/5 + 7 1/3 ] . In your book, are their parentheses around the 2/3 and the 1/3 ?

OpenStudy (anonymous):

nope

Directrix (directrix):

Okay. Now worries but I need to recalculate. What is 6 2/5 as an improper (top heavy) fraction? Ditto for 7 1/3.

Directrix (directrix):

2 [ 6 2/5 + 7 1/3 ] = 2 [ 32/5 + 22/3 ] = (agree so far?)

OpenStudy (anonymous):

yes

Directrix (directrix):

2 [ 32/5 + 22/3 ] = Distributive property time so 2 [ 32/5 ] + 2 [ 22/3] =

Directrix (directrix):

2 [ 32/5 ] + 2 [ 22/3] = 64/5 + 44/3 = ---> Agree?

Directrix (directrix):

192/15 + 220/15 = 412/15 -- improper fraction = 9 7/15 as mixed number (no parenthesis) --------> I think this is the one. Check. = 27.47 approx

Directrix (directrix):

9 7/15 as in 9 and seven-fifteenths

Directrix (directrix):

Strange that this rounds to 9.5 but you were doing mixed numbers from the start.